OpenStudy (anonymous):

Choose the correct simplification of (3x4y5)2(y2)3.

OpenStudy (anonymous):

Where to choose from?

OpenStudy (anonymous):

\[\LARGE (3x^4y^5)^2(y^2)^3\] If we're talking about this then ... here is how it goes. use \(\LARGE (x^a)^b=x^{a\cdot b}\) and you have... \[\LARGE (3x^4y^5)^2(y^2)^3=(9x^8y^{10})(y^6)=\] now using \(\LARGE x^a\cdot x^b=x^{a+b}\) you'd have... \[\LARGE 9x^8y^{10+6}=9x^8y^{16}\]

OpenStudy (anonymous):

I actually didn't mean to post that question so thank you very much for your reply's ( : It's greatly appreciated.